tvmara:

Factor the following complex polynomials using the decomposition method: show all work. a. 15x^2 + x - 6

XxAngelGirlxX:

Factoring 15x2-x-6 The first term is, 15x2 its coefficient is 15. The middle term is, -x its coefficient is -1. The last term, "the constant", is -6 Step-1: Multiply the coefficient of the first term by the constant 15 • -6 = -90 Step-2: Find two factors of -90 whose sum equals the coefficient of the middle term, which is -1. -90 + 1 = -89 -45 + 2 = -43 -30 + 3 = -27 -18 + 5 = -13 -15 + 6 = -9 -10 + 9 = -1 That's it Step-3: Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -10 and 9 15x2 - 10x + 9x - 6 Step-4 : Add up the first 2 terms, pulling out like factors : 5x • (3x-2) Add up the last 2 terms, pulling out common factors : 3 • (3x-2) Step-5: Add up the four terms of step 4 : (5x+3) • (3x-2) Which is the desired factorization Final Answer - (3x - 2) • (5x + 3)